The production function of given firm Is given by Q =200\sqrtLK  and the.

associated cost function is given by 2L + 5K = 1000

A. Calculate the level of labor and capital input at the optimum level of production

B. Calculate the MRTSL,K and MRTSk, L

c. What is the level of output at the least cost combination of labor and capital inputs?

1
Expert's answer
2021-09-23T09:13:41-0400
"MU_K=100\sqrt{\frac{L}{K}}"

"MU_L=100\sqrt{\frac{K}{L}}"

"K=100"

"L=250"

"MRTS_{LK}=\frac{250}{100}=2.5"

"MRTS_{KL}=\frac{100}{250}=0.4"

"Q=100\sqrt{100\times250}=15811.4"